The Wage is $40. The Land costs $100. The wheat sells for $10 per unit.

1. The MARGINAL product of the 5th worker is equal to?

2. The MARGINAL product of the 7th worker is equal to?

3. With 7 workers TOTAL COST is equal to?

4. With 4 workers TOTAL REVENUE is equal to?

5. With 7 workers TOTAL REVENUE is equal to?

6. If all these workers have the same skill level, why does worker 2 have a LARGER Marginal product than worker 1?

7. How many workers should this firm hire to maximize its profits?

8. If this firm does hire the number of workers you answered for questions (7), the will have TOTAL PROFITS equal to?

9. If all these workers have the same level of skill, why does worker 6 have a smaller marginal product than worker 5?

10. What is the a mount of TOTAL Fixed costs for this firm in the SHORT run?

#of worker Total Product

1 ..........................50

2 .........................150

3. ........................300

4 .........................350

5 ........................360

6 .......................365

7 ........................369
Answer #1

(1)

Marginal product (MP) = Change in total product (Q) / Change in labor (L)

When L = 4, Q = 350 and when L = 5, Q = 360.

MPL=5 = (360 - 350) / (5 - 4) = 10/1 = 10

(2)

When L = 6, Q = 365 and when L = 7, Q = 369.

MPL=5 = (369 - 365) / (7 - 6) = 4/1 = 4

(3)

Total cost ($) = wL + Fixed cost = 40 x 7 + 100 = 280 + 100 = 380

(4)

Total revenue ($) = P x Q = 10 x 350 = 3500

(5)

Total revenue ($) = 10 x 369 = 3690

NOTE: As per Answering Policy, 1st 5 parts are answered.